OpenStudy (anonymous):

y=1/3x-5 in slope intercept form

OpenStudy (acxbox22):

y=mx+b is slope intercept form where m is slope and b is y intercept acording to your equation it is already in slope intercept form

OpenStudy (anonymous):

thanks

OpenStudy (acxbox22):

;)

OpenStudy (anonymous):

how am i suppose to graph it

OpenStudy (anonymous):

-5 is the y intercept, so put a point on -5 on the y axis

OpenStudy (anonymous):

1/3= rise/run go up one unit from -5 and 3 across to the right

OpenStudy (anonymous):

so wat r the points